# LocalTweets to LocalHealth: A Mental Health Surveillance Framework Based on Twitter Data

### v1

@@ Abstract
Prior research on Twitter (now X) data has provided positive evidence of its utility in developing supplementary health surveillance systems. In this study, we present a new framework to surveil public health, focusing on mental health (MH) outcomes. We hypothesize that locally posted tweets are indicative of local MH outcomes and collect tweets posted from 765 neighborhoods (census block groups) in the USA. We pair these tweets from each neighborhood with the corresponding MH outcome reported by the Center for Disease Control (CDC) to create a benchmark dataset, LocalTweets. With LocalTweets, we present the first population-level evaluation task for Twitter-based MH surveillance systems. We then develop an efficient and effective method, LocalHealth, for predicting MH outcomes based on LocalTweets. When used with GPT3.5, LocalHealth achieves the highest F1-score and accuracy of 0.7429 and 79.78%, respectively, a 59% improvement in F1-score over the GPT3.5 in zero-shot setting. We also utilize LocalHealth to extrapolate CDC’s estimates to proxy unreported neighborhoods, achieving an F1-score of 0.7291. Our work suggests that Twitter data can be effectively leveraged to simulate neighborhood-level MH outcomes.thanks: Data and scripts will be linked soon. Corresponding Author: vijeta_deshpande@student.uml.edu
@@
[…]
${{}^{1}}$niversity of Massachusetts Lowell,
${{}^{2}}$niversity of Minnesota,
${{}^{3}}$niversity of Massachusetts Amherst,
${{}^{4}}$niversity of Colorado Anschutz Medical Campus,
${{}^{5}}$niversity of Massachusetts, Chan Medical School,
@@ 2.   Related Works
Figure 2: Distributional Properties of LocalTweets. (A,B) MH tweets are the most numerous, while FI and general tweets have comparable volumes. Tweet volume is slightly skewed toward the South and West regions and toward ADIs $\geq 70$ and $\leq 20$. (C,D) The distribution of data splits over regions and ADIs is approximately the same. The distribution of all BGs over ADI values is also approximately balanced.
@@ 3.4.   Cleaning and splitting the data for model development
We filter the Twitter-CDC coupled data to remove BGs with no tweets (for any category and any year), and BGs not included in the CDC data. This filtering process removed 235 BGs, resulting in a final dataset of 765 BGs with corresponding Twitter and CDC data for five years. We refer to the cleaned version of the data as LocalTweets. Figure 2 shows the distributional properties of the LocalTweets. We split the data differently for our two experimental settings namely, the forecasting setting and spatial extrapolation setting. For forecasting, we use all 2019 data as the test dataset and consider 2015 to 2018 data for model development. For the spatial extrapolation, we first divide the final set of 765 BGs into three splits (test, train, and validation) such that the distribution over ADI values and geographic regions is approximately held the same across splits, refer to Figure 2, panels C and D. We remove the BGs in test split, to create a proxy set of "unreported" BGs. We use 2015-19 data from the train and validation split for model development and report the performance on the unreported BGs (test split) of the year 2019.
@@
where the subscripts b and y represent the BG and year, the superscripts k and s denote the tweet category and split. The variables t, c, and g represent the tweets, tweet counts, and ground truth outcome values, respectively. We create a risk category variable $r$, such that $r^{(s)}_{b,y}=1$ (high-risk BG) if $g^{(s)}_{b,y}\geq 75^{th}$ percentile, otherwise $r^{(s)}_{b,y}=0$ (low-risk BG). We use the variable $r$ to evaluate various models. Note that the superscript k is not present for $g$ and $r$ variables because we only consider the MH outcome as the target. In Appendix B, we provide additional distributional details of LocalTweets. Furthermore, in Appendix C, we show fewer BGs in the Northeast region do not affect the geographic generalizability of LocalTweets.
@@ 4.   Methodology
Our methods of analysis are mainly divided into two parts. First, we conduct a correlation analysis to investigate the agreement between Twitter data-based statistics and the reported CDC outcome values. In the second part, we present a regression analysis. Specifically, we develop a model that processes the Twitter data (set of tweets) to predict the MH outcome value for respective neighborhoods, refer to Appendix D for a simple schematic. In the following subsections, we discuss each part in detail.
@@ 4.2.   Regression analysis
Such that, $|t_{b_{s},y}^{(k)}|=\min(4000,|t_{b,y}^{(k)}|)$ $\forall\ b,y,k$ Step : Encoding
@@
Such that, $v_{b_{s},y}^{(i)}=[v_{b_{s},y,1}^{(k)},v_{b_{s},y,2}^{(k)},\ldots,v_{b_{s},y,n}^{(k)}]$, where $v_{b_{s},y,j}^{(k)}$ is the representation vector of the $j^{th}$ tweet in the $t_{b_{s},y}^{(k)}$ set and $n=|t_{b_{s},y}^{(k)}|$ Step : Aggregation
@@ 5.1.   Sets of experiments
Set-2: Effect of text encoder. In this experiment, we replace the language model ($f_{LM}(\cdot)$) in Algorithm 1 with various pre-trained language models, including Twitter-RoBERTa Barbieri et al. (2020), PHS-BERTNaseem et al. (2022), and GPT-3.5 OpenAI (2021), and assess the changes on forecasting performance. We work with “general” category of tweets and consider 2015 to 2018 data for developing the model and, 2019 data for testing. We also measure the zero-shot performance of GPT-3.5, refer to Appendix E for further details.
@@ 5.4.   Hyperparameters and evaluation
We train all models for 1,600 epochs with a batch size of 512. We use a linear learning rate schedule with a 20% warmup and peak learning rate of $1\times 10^{-3}$. We minimize mean squared error (MSE) using AdamW Loshchilov and Hutter (2017); PyTorch (2023) with a weight decay of 0.1. We evaluate models with macro-F1 in identifying the risk category ($r$) of BGs and save the best model based on validation macro-F1. We evaluate the best model on the test split and report averaged macro-F1 and accuracy across 10 random seeds. For both, SVM and LoR, we change the loss function to binary cross-entropy and we use a classification threshold of 0.15 to identify high-risk BGs.
@@ 6.   Results
MH only
0.5668 60.05
FI only
0.5602 64.43
MH and FI
0.5853 64.16
General only
0.5984 66.76
@@
MH only
0.7089 74.52
FI only
0.7117 75.36
MH and FI
0.7085 74.33
General only
0.7236 76.48
@@
For text-based models, a comparison within the tweet categories revealed interesting insights. The model with general tweets performed better than other categories of tweets. This finding supports our hypothesis (refer to Section 3.2) and highlights the better generalization capabilities of the general tweets for population-level MH outcome prediction, compared to the keyword-derived tweets. Hence, our finding motivates the usage of general tweets for the prediction of population-level MH outcomes.
@@
Majority baseline
0.4336 76.56
GPT3.5 (0-shot)
0 0.4675 76.21
ADI only
2 0.6406 72.81
RoBERTa-base
210 0.7236 76.48
RoBERTa-large
274 0.7228 76.04
Twitter-RoBERTa-base
210 0.7245 76.44
PHS-BERT
274 0.7301 76.97
GPT3.5
402 0.7429 79.78
@@
Contrary to our expectations, we observed a slight declination in the F1-score when we augmented the 2015 data with the data from 2016 to 2018 for developing the model, refer to Figure 3. The declination is 0.0036 for GPT3.5 and 0.0127 for RoBERTa-base. This unexpected dip may be attributable to the possibility that Twitter posts in 2015 may not accurately represent the population MH status in 2019. Interestingly, we found that RoBERTa (with an F1-score of 0.7042) outperforms GPT-3.5 (F1-score: 0.6406) when only 2018 data is available to develop the model. In other cases, when more data is available, GPT-3.5 presents as a better choice of tweet encoder.
@@ 9.   Ethics Statement
While this study demonstrates the potential utility of Twitter data for supplementary mental health surveillance, we acknowledge important ethical considerations. First, the use of social media data raises privacy concerns. We took steps to protect privacy by only collecting tweets that were publicly accessible, aggregating tweets at the neighborhood level, and not analyzing individual users’ data. Second, while supplementary surveillance has benefits, it cannot capture lived experiences. Thus findings should be considered preliminary and complemented by qualitative, participatory research methods. Third, mental health is a sensitive issue, and care must be taken not to further stigmatize mental illness. While our work aims for the betterment of mental health public policies, we acknowledge that the findings of our study could be used to develop algorithms that can target distressed areas or populations at risk with discriminatory or harmful content. Lastly, on the technical front, we made our best efforts to reduce the technical barrier to research by considering economic language models, and training lean ($\leq 402$ parameters) systems. Our goal was to encourage community participation for the benefit of the community. However, we recognize that the barrier is also contingent upon Twitter’s data-sharing and pricing policies.
@@ Appendix A Collection of Tweets
Mental health
’bored’, ’disgusting’, ’sick of’, ’tired of it’, ’dont want to’, ’so fucking miserable’, ’tired of being’, ’depressed’, ’alone’, ’isolate’, ’given up’, ’no friend’, ’cant deal’, ’want to talk’, ’in my room’, ’awake’, ’sleepless’, ’nightmares’, ’insomnia’, ’cant sleep’, ’wish sleep’, ’up all night’, ’body is begging’, ’exhausted’, ’tired’, ’my energy’, ’dont have energy’, ’tired to look’, ’feel myself falling’, ’binge’, ’fasting’, ’eating disorder’, ’eat again’, ’always eating’, ’forced to eat’, ’am eating ?’, ’failure’, ’ugly’, ’worthless’, ’hate myself’, ’fat piece’, ’self hatred’, ’piece of shit’, ’feel like trash’, ’thoughts’, ’confused’, ’overthinking’, ’am losing’, ’losing mind’, ’my mind off’, ’quiet’, ’attention’, ’nervous’, ’social anxiety’, ’dead quiet’, ’dont wanna move’, ’cut’, ’hang’, ’blade’, ’die’, ’suicidal’, ’rip skin’, ’suicide attempt’, ’car hit’, ’kill myself’, ’of the road’
Food insecurity
"food stamps", "SNAP", "food charities", "food pantry", "food voucher", "deficiency", "hunger", "hungry", "food insecurity", "poor diet", "junk food", "food desert", "poor nutrition", "starvation", "without food", "no food", "no groceries", "lack of food", "not enough food"
General
@@ Appendix B Data Properties
Appendix C Effect of Skewed Regional Distribution

### v2

@@ Abstract
Prior research on Twitter (now X) data has provided positive evidence of its utility in developing supplementary health surveillance systems. In this study, we present a new framework to surveil public health, focusing on mental health (MH) outcomes. We hypothesize that locally posted tweets are indicative of local MH outcomes and collect tweets posted from 765 neighborhoods (census block groups) in the USA. We pair these tweets from each neighborhood with the corresponding MH outcome reported by the Center for Disease Control (CDC) to create a benchmark dataset, LocalTweets. With LocalTweets, we present the first population-level evaluation task for Twitter-based MH surveillance systems. We then develop an efficient and effective method, LocalHealth, for predicting MH outcomes based on LocalTweets. When used with GPT3.5, LocalHealth achieves the highest F1-score and accuracy of 0.7429 and 79.78%, respectively, a 59% improvement in F1-score over the GPT3.5 in zero-shot setting. We also utilize LocalHealth to extrapolate CDC’s estimates to proxy unreported neighborhoods, achieving an F1-score of 0.7291. Our work suggests that Twitter data can be effectively leveraged to simulate neighborhood-level MH outcomes.thanks: Following responsible data practices, we will share data for requests that align with our privacy policy. Corresponding Author: vijeta_deshpande@student.uml.edu
@@
Vijeta Deshpande1, Minhwa Lee2, Zonghai Yao3, Zihao Zhang3, Jason Brian Gibbons4, Hong Yu1,3,5
1University of Massachusetts Lowell,
2University of Minnesota,
3University of Massachusetts Amherst,
4University of Colorado Anschutz Medical Campus,
5University of Massachusetts, Chan Medical School,
@@ 2.   Related Works
Figure 2: Distributional Properties of LocalTweets. (A): Region vs. Number of Tweets: MH tweets are the most numerous, while FI and general tweets have comparable volumes. Tweet volume is slightly skewed toward the South and West regions. (B) ADI vs. Number of Tweets: MH and FI tweets are slightly skewed toward ADIs $\geq 70$ and $\leq 20$. (C) Region vs. Number of BGs: The distribution of data splits over regions is approximately the same. The number of BGs from the Northeast region is less than other regions. Refer to Appendix E for more discussion. (D): ADI vs. Number of BGs: The number of BGs is fairly balanced over the ADI values and across the data splits.
@@ 3.4.   Cleaning and splitting the data for model development
We filter the Twitter-CDC coupled data to remove BGs with no tweets (for any category and any year), and BGs not included in the CDC data. This filtering process removed 235 BGs, resulting in a final dataset of 765 BGs with corresponding Twitter and CDC data for five years. We refer to the cleaned version of the data as LocalTweets. Figure 2 shows the distributional properties of the LocalTweets. We provide a detailed description of data properties in Appendices B and C. In the cleaned version of LocalTweets, we observe significantly fewer BGs from the Northeast region. Hence, we analyzed the effect of fewer BGs from the Northeast region on the space generalizability of the data. We observe that fewer BGs from the Northeast region do not hamper the space generalizability of the data and provide details in Appendix E.
We split the data differently for our two experimental settings namely, the forecasting setting and spatial extrapolation setting. For forecasting, we use all 2019 data as the test dataset and consider 2015 to 2018 data for model development. For the spatial extrapolation, we first divide the final set of 765 BGs into three splits (test, train, and validation) such that the distribution over ADI values and geographic regions is approximately held the same across splits, refer to Figure 2, panels C and D. We remove the BGs in test split, to create a proxy set of "unreported" BGs. We use 2015-19 data from the train and validation split for model development and report the performance on the unreported BGs (test split) of the year 2019.
@@
where the subscripts b and y represent the BG and year, the superscripts k and s denote the tweet category and split. The variables t, c, and g represent the tweets, tweet counts, and ground truth outcome values, respectively. We create a risk category variable $r$, such that $r^{(s)}_{b,y}=1$ (high-risk BG) if $g^{(s)}_{b,y}\geq 75^{th}$ percentile, otherwise $r^{(s)}_{b,y}=0$ (low-risk BG). In other words, $r$ is nothing but a flag indicating whether a BG is high-risk or not. We leverage the variable $r$ to evaluate regression models’ performance with discrete metrics (e.g., accuracy and F1-score). Note that the superscript k is not present for $g$ and $r$ variables because we only consider the MH outcome as the target variable.
@@ 4.   Methodology
Our methods of analysis are mainly divided into two parts. First, we conduct a correlation analysis to investigate the agreement between Twitter data-based statistics and the reported CDC outcome values. In the second part, we present a regression analysis. Specifically, we develop a model that processes the Twitter data (set of tweets) to predict the MH outcome value for respective neighborhoods, refer to Appendix G for a simple schematic. In the following subsections, we discuss each part in detail.
@@ 4.2.   Regression analysis
Such that, $|t_{b_{s},y}^{(k)}|=\min(4000,|t_{b,y}^{(k)}|)$ $\forall\ b,y,k$
Step : Encoding
@@
Such that, $v_{b_{s},y}^{(i)}=[v_{b_{s},y,1}^{(k)},v_{b_{s},y,2}^{(k)},\ldots,v_{b_{s},y,n}^{(k)}]$, where $v_{b_{s},y,j}^{(k)}$ is the representation vector of the $j^{th}$ tweet in the $t_{b_{s},y}^{(k)}$ set and $n=|t_{b_{s},y}^{(k)}|$
Step : Aggregation
@@ 5.1.   Sets of experiments
Set-2: Effect of text encoder. In this experiment, we replace the language model ($f_{LM}(\cdot)$) in Algorithm 1 with various pre-trained language models, including Twitter-RoBERTa Barbieri et al. (2020), PHS-BERTNaseem et al. (2022), and GPT-3.5 OpenAI (2021), and assess the changes on forecasting performance. We work with “general” category of tweets and consider 2015 to 2018 data for developing the model and, 2019 data for testing. We also measure the zero-shot performance of GPT-3.5, refer to Appendix F for further details.
@@ 5.4.   Hyperparameters and evaluation
We train all models for 1,600 epochs with a batch size of 512. We use a linear learning rate schedule with a 20% warmup and peak learning rate of $1\times 10^{-3}$. We minimize mean squared error (MSE) using AdamW Loshchilov and Hutter (2017); PyTorch (2023) with a weight decay of 0.1. In order to leverage standard classification metrics for model evaluation, we employ a thresholding technique to convert the continuous-valued model outputs ($\hat{g}_{b,y}$) into binary risk category predictions. This allows us to directly compare predicted risk categories with the ground truth labels (variable $r$, Section 3.4) using established metrics like accuracy and F1-score (macro-averaged). Model selection is conducted based on the macro-F1 score achieved on the validation set. Following training, we evaluate the best model on the test split and report averaged macro-F1 and accuracy across 10 random seeds. For both, SVM and LoR, we change the loss function to binary cross-entropy and we use a classification threshold of 0.15 to identify high-risk BGs.
@@ 6.   Results
MH only 0.5668 60.05
FI only 0.5602 64.43
MH and FI 0.5853 64.16
General only 0.5984 66.76
@@
MH only 0.7089 74.52
FI only 0.7117 75.36
MH and FI 0.7085 74.33
General only 0.7236 76.48
@@
For text-based models, a comparison within the tweet categories revealed interesting insights. The model with general tweets performed better than other categories of tweets. This finding supports our hypothesis (refer to Section 3.2) and highlights the better generalization capabilities of the general tweets for population-level MH outcome prediction, compared to the keyword-derived tweets. Hence, our finding motivates the usage of general tweets for the prediction of population-level MH outcomes. We provide additional comparison of the statistical properties of tweet categories in Appendix D.
@@
Majority baseline 0.4336 76.56
GPT3.5 (0-shot) 0 0.4675 76.21
ADI only 2 0.6406 72.81
RoBERTa-base 210 0.7236 76.48
RoBERTa-large 274 0.7228 76.04
Twitter-RoBERTa-base 210 0.7245 76.44
PHS-BERT 274 0.7301 76.97
GPT3.5 402 0.7429 79.78
@@
Contrary to our expectations, we observed a slight declination in the F1-score when we augmented the 2015 data with the data from 2016 to 2018 for developing the model, refer to Figure 3. The declination is 0.0036 for GPT3.5 and 0.0127 for RoBERTa-base. This unexpected dip may be attributable to the possibility that Twitter posts in 2015 may not accurately represent the population MH status in 2019. Interestingly, we found that RoBERTa (with an F1-score of 0.7042) outperforms GPT-3.5 (F1-score: 0.6406) when only 2018 data is available to develop the model. In other cases, when more data is available, GPT-3.5 presents as a better choice of tweet encoder. To this end, we compared the statistical properties of RoBERTa-base and GPT-3.5. We observe a higher variance in the encodings taken from the RoBERTa model and speculate that the more spread out data potentially helps LocalHealth model to detect underlying patterns for MH prediction (refer to Appendix C).
@@ 9.   Ethics Statement
While this study demonstrates the potential utility of Twitter data for supplementary mental health surveillance, we acknowledge important ethical considerations. In this section, we describe the procedure we adopted to ensure rightful data access, privacy preservation, and gated sharing of the data.
Data Access: To access Twitter data we followed the Twitter Developer Account application procedure [^5] Updated procedure and terms and conditions can be found HERE.. Our application for accessing Twitter data was reviewed, scrutinized, and approved by Twitter, based on an academic research proposal focused on leveraging Twitter data for public health applications.
Data Privacy Preservation: The use of social media data raises privacy concerns. We took rigorous steps throughout our analysis to protect the privacy of the data. Firstly, we selected tweets that are publicly available and did not collect data from any profiles or tweets that are marked private. For model development, we focused on tweet texts only and did not make use of any additional features of the tweet or the user such as, location or demographic features. We used twarc library to fetch tweets posted from a specific location (BGs) but we did not access or utilize the location feature of the tweets for model development. Furthermore, in the presented study we ensured privacy preservation through encoding and aggregation of the tweets. In the first step textual information gets encoded into high-dimensional vectors and thousands of such vectors are aggregated together for each block-group. With these two steps, we ensure that human-readable text cannot be excavated from the aggregated representation.
Data Bias: The final version of LocalTweets likely is biased along demographic characteristics. However, to maintain the privacy of users and ethical usage of the collected data we did not explore bias along demographic features. We assume that the bias exists and we improve the performance of LocalHealth within the constraints of the demographic bias. Nonetheless, we maintain a fairly balanced distribution across ADI values. We clearly show the existing regional bias and conduct detailed analysis to show that it does not affect results majorly. Furthermore, we show that the bias derived from the commonly used keyword-based data collection methods does not generalize well. Hence, we make the best effort to address bias-related issues while maintaining privacy of the data.
Usage of LocalHealth: We study the usage of LocalTweets and the application of LocalHealth solely as a supplementary system for traditional health surveillance systems. While supplementary surveillance has benefits, it cannot capture lived experiences. Thus our findings should be considered preliminary and complemented by qualitative, participatory research methods. In addition, mental health is a sensitive topic, and care must be taken not to further stigmatize mental illness. While our work aims for the betterment of mental health public policies, we acknowledge that the findings of our study could be used to develop algorithms that can target distressed areas or populations at risk with discriminatory or harmful content. Hence, we will provide gated access to LocalHealth. We will release the model and the data based on individual requests that adhere to, (1) focus usage of the data for research on public health research questions (2) follow Twitter’s data privacy policy.
Reproducability: Lastly, on the technical front, we made our best efforts to reduce the technical barrier to research by considering economic language models, and training lean ($\leq 402$ parameters) systems. Our goal was to encourage community participation for the benefit of the community. However, we recognize that the barrier is also contingent upon Twitter’s privacy policies.
@@ Appendix A Collection of Tweets
Mental health ’bored’, ’disgusting’, ’sick of’, ’tired of it’, ’dont want to’, ’so fucking miserable’, ’tired of being’, ’depressed’, ’alone’, ’isolate’, ’given up’, ’no friend’, ’cant deal’, ’want to talk’, ’in my room’, ’awake’, ’sleepless’, ’nightmares’, ’insomnia’, ’cant sleep’, ’wish sleep’, ’up all night’, ’body is begging’, ’exhausted’, ’tired’, ’my energy’, ’dont have energy’, ’tired to look’, ’feel myself falling’, ’binge’, ’fasting’, ’eating disorder’, ’eat again’, ’always eating’, ’forced to eat’, ’am eating ?’, ’failure’, ’ugly’, ’worthless’, ’hate myself’, ’fat piece’, ’self hatred’, ’piece of shit’, ’feel like trash’, ’thoughts’, ’confused’, ’overthinking’, ’am losing’, ’losing mind’, ’my mind off’, ’quiet’, ’attention’, ’nervous’, ’social anxiety’, ’dead quiet’, ’dont wanna move’, ’cut’, ’hang’, ’blade’, ’die’, ’suicidal’, ’rip skin’, ’suicide attempt’, ’car hit’, ’kill myself’, ’of the road’
Food insecurity "food stamps", "SNAP", "food charities", "food pantry", "food voucher", "deficiency", "hunger", "hungry", "food insecurity", "poor diet", "junk food", "food desert", "poor nutrition", "starvation", "without food", "no food", "no groceries", "lack of food", "not enough food"
General " "
@@
Figure 5: Longitudinal Trend of Target Variable. In this figure, we present the average values of MH outcomes for the years 2015 to 2019. The average value is calculated separately for each ADI value considered in the analysis. The increasing trend in the MH outcome values is observed across all BGs irrespective of their socio-economic status.
Figure 6: Average of Activations. In this figure we present the average value of the last hidden state for RoBERTa and GPT-3.5 model. The average values are calculated separately for each year, over all dimensions of the hidden vector and 765 block-groups. We observe fairly stable average values, primarily due to the normalization operations included in the language models.
Figure 7: Standard Deviation of Activations. In this figure we present the standard deviation (STD) of the last hidden state for RoBERTa and GPT-3.5 model. The STD values are calculated separately for each year, the overall dimensions of the hidden vector, and 765 block-groups. We observe a significant difference between the STD values for RoBERTa and GPT-3.5. In addition, we also note a slight increase in STD for RoBERTa in the year 2018 and 2019.
@@ Appendix B Data Properties
Figure 8: Average Across Tweet Categories. In this figure we present the average of the RoBERTa last hidden state for all tweet categories. The average values are calculated separately for each year, the overall dimensions of the hidden vector, and 765 block-groups. We observe negligible differences between tweet categories.
Appendix C Longitudinal Properties of Input and Target Variables
In this section, we present a few longitudinal properties of the input and target variables. To reiterate, the input variable is the encoding of a set of tweets from a specific language model (refer to Section 4.2). The target variable is the MH outcomes collected from the CDC database.
In Figure 5, we present the variation of the MH outcomes value in time. We calculate average values of MH outcomes across all BGs (765), separately for each year. We observe that MH outcomes have a consistent increasing trend over the years. Interestingly, this trend holds irrespective of the socio-economic status (ADI) of BGs. Hence, to effectively solve the problem of predicting MH outcomes, the model needs to recognize patterns in the tweets that eventually lead to an increasing pattern in the MH outcomes.
The input variables i.e., the encoding from language model and hence, high dimensional. Hence, to briefly understand the longitudinal patterns in the input variables, we calculate mean and standard deviation values, over all encoding dimensions and BGs, but separately for each year, refer to Figures 6 and 7. For both language models, the mean values are stable for all years. The mean values for GPT-3.5 are much closer to zero as compared to the RoBERTa model. The reason for such robust mean values is the normalization operations conducted in the language models. While mean values for both models are close to each other, there are noticeable differences in the standard deviation values. The standard deviation of the RoBERTa model is approximately 16 times higher than GPT-3.5. In addition, there also exists a slight increase in the standard deviation value in the RoBERTa model’s activations for the year 2018 and 2019. Such increment in the standard deviation values is not observed in the GPT-3.5 model. We believe these statistical properties are the primary reason behind the performance differences between RoBERTa and GPT-3.5 models presented in Figures 3 and 4. The high variance of the RoBERTa encodings possibly helps the model to identify underlying hidden patterns effectively.
Figure 9: Standard Deviation Across Tweet Categories. In this figure we present the standard deviation (STD) of the RoBERTa last hidden state for all tweet categories. The STD values are calculated separately for each year, the overall dimensions of the hidden vector, and 765 block-groups. We observe a minor difference between tweet categories. Notably, the variance in the general category tweets is consistently observed as the lowest among all categories, for all years.
Appendix D Properties of Tweet Categories
Our experiment results presented in Table 2 highlight the advantage of using general category tweets over the other mental health (MH) associated categories namely, food insecurity (FI) and MH. Here, we present the mean and standard deviation of RoBERTa encodings for all three categories. In the LocalTweets dataset, we notice only minor differences between the mean and standard deviation for various tweet categories, refer to Figure 8, 9. Nonetheless, we observe the standard deviation value for general category tweets to be the lowest, consistently over all years. We speculate that the lower variance in the general category tweets might help focus the model on a specific semantic latent space that is potentially beneficial for the prediction of MH outcomes.
Appendix E Effect of Skewed Regional Distribution